find each unit price to decide which is the best buy.
[A] 6 hot dog for $6.06
[B]16 hot dogs for $16.32
[c] 10 hot dogs for $9.90
[d] 8 hot dogs for $8.32

Answer by bamaboy(14)

find each unit price to decide which is the best buy.
[A] 6 hot dog for $6.06
6.06/6 = 1.01
[B]16 hot dogs for $16.32
16.32 / 16 = 1.02
[c] 10 hot dogs for $9.90
9.90 / 10 = .99
[d] 8 hot dogs for $8.32
8.32 / 8 = 1.04
so b is your best bet because u can get 6 more hot dogs for only .03 cent